A 1.5ton automobile crashed into a wall at a speed of 10mi/h. the car moved 5in. before being

brought to rest. What was the average force exerted on the wall be the car?


1
Expert's answer
2022-03-29T16:46:23-0400

Given:

"m=1500\:\rm kg"

"v_0=10\:\rm mi/h=4.47\:\rm m/s"

"d=5\:\rm in=0.127\: m"


The average force exerted on the wall by a car

"F=ma=\frac{mv_0^2}{2d}=\frac{1500*4.47^2}{2*0.127}=1.18*10^5\:\rm N"